40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8808     Entry time: Tue Jul 9 01:18:48 2013
Author: Nic, Koji 
Type: Update 
Category: ASC 
Subject: PRMI locking / PRM ASC adjustment 

[Koji, Nic]

- Locked PRMI with REFL165 I/Q

- Aligned the POP beam on the QPD. We found that the vertical motion of the beam appeared in the yaw signal, and horizontal motion in the pitch signal.
  This was fixed by swapping the cables to the ADC. Later it turned out that this was caused by the calibration setup for the QPD.
  We requested Jenne to fix the QPD on the table with the current orientation.

- Re-implemented the AC-coupled ASC servo. The filters were just copied from the previous PRM ASC servo (in the SUS ASC filter).
  The same filter was installed to the pitch and yaw filter modules for now. The gains were adjusted to have some stable lock stretches.
  C1:ASC-PRCL_YAW_GAIN: -0.01
  C1:ASC-PRCL_PIT_GAIN: -0.01

  The power spectra of C1:ASC-PRCL_YAW_IN1 and C1:ASC-PRCL_PIT_IN1 were attached.
  The reference curves are the ones with the servo on. The other two are the free-running stability of the QPD output.

- Modified the up and down scripts for the PRM ASC for the new setup.
  It first turns on the inputs of the filters and then turn on FM2/3.
  It assumes that the outputs are engaged all time.

 

Attachment 1: PRMI_ASC.pdf  47 kB  Uploaded Tue Jul 9 02:19:26 2013  | Hide | Hide all
PRMI_ASC.pdf
ELOG V3.1.3-